# LiT: Delving into a Simplified Linear Diffusion Transformer for Image Generation

### v1

@@ 4.3 Inheriting Weights from Diffusion Transformer
4.4 Distilling Both Predicted Noise and Variance
A pre-trained DiT-S/2/XL/2 serves as the teacher network to distill the student LiT-S/2. For strong performance, we use a DiT-S/2 pre-trained for 800K iterations for initialization, following Guideline 3&4. Denote as the noised input latent in timestep , where , , are height, width and the number of channels of the input latent. denote network parameters of the teacher diffusion Transformer and the student LiT network , respectively. As denoising diffusion models, the output of DiT and LiT include both noise predictions and variances of the reverse diffusion process . DDPM [29] fixes the variance and optimizes model parameters using . Naturally, a straightforward idea is to align the student network’s noise prediction capabilities with those of the teacher through a mean squared error (MSE) objective. Thus, we have:
where is the normalization factor. The impact of distilling the predicted noise is reported in Tab. 4. It shows that using a stronger teacher model (e.g., DiT-XL/2) improves the student model’s generation capbility (51.13 vs. 53.83, measured by FID). Moreover, increasing the weight of the distillation objective from 0.05 to 0.5 consistently boosts performance, suggesting that the cheap linear attention can benefit from the supervision of the self-attention in diffusion Transformers.
@@ 4.4 Distilling Both Predicted Noise and Variance
As demonstrated in IDDPM [52], a suitable variance can contribute to the initial steps of the diffusion process, thereby improving the log-likelihood. Accordingly, the authors set the variance learnable. Based on the finding, we propose to use the variance from the teacher diffusion model as a supervision for the student, but in a moderate way. The MSE between the reverse process variances of teacher and student can be calculated as:
Closing remarks.
So far, we have finalized the roadmap. LiT follows the macro/micro-level design of DiT but employs efficient linear attention. Using our cost-effective training strategies, LiT-S/2 significantly improves the FID. Next, we validate it on larger variants (e.g., B/L/XL) and challenging tasks (e.g., text-to-image generation).
5 Experiments
5.1 Class-Conditional Image Generation
256256 ImageNet.
We first evaluate LiT on the 256256 ImageNet [13] benchmark for conditional image generation. The LiT-S/2, B/2, L/2, XL/2 configurations are consistent with DiT, except that the linear attention heads are set to 2/3/4/4, following the guideline 2 in Sec. 4.2. DWC kernel size is set to 5 for all model variants. Detailed configuration are provided in Sec. A of the appendix. We train 400K steps with a batch size of 256. For LiT-XL/2, we extend training to 1.4M steps (20 of DiT’s [53] 7M steps). We used the AdamW [47] optimizer with a learning rate of and weight decay set to 0. Following Sec. 4.3, we initialize LiT’s parameters using the pre-trained DiT. , in hybrid knowledge distillation, are set to 0.5 and 0.05. Evaluation metrics include FID-50K [27], Inception Score [65] and Precision/Recall [40].
Tab. 5 compares the results of different model variants of LiT and DiT. Notably, LiT trained for only 100K steps already outperforms the DiT trained for 400K steps across various evaluation metrics and different model variants. With extra training up to 400K steps, the model’s performance continues to improve. In Tab. 6, we compare LiT-XL/2 trained for 1.4M steps with high-performance baselines. Despite having only of the training steps of DiT-XL/2, LiT still competes on par with DiT (2.32 vs. 2.27). Besides, LiT competes favorably with several U-Net [63] based baselines. The results demonstrate that the linear attention, when combined with suitable optimization strategies, can be reliably used in image generation applications.
@@ 5.1 Class-Conditional Image Generation
We further validate LiT-XL/2 on the 512512 ImageNet benchmark. Using the pre-trained DiT-XL/2 [53] as the teacher model, we initialize LiT-XL/2 with its weights. For knowledge distillation, we set and to 1.0 and 0.05, respectively, and train LiT-XL/2 for only 700K steps ( of DiT’s [53] 3M steps). Notably, unlike DiT, which used a batch size of 256, we adopt a smaller batch size of 128, requiring twice the training steps to complete 1 epoch (i.e., processing the entire training dataset once). Aside from this, all other hyper-parameter settings were consistent with the 256256 experiments. The results are reported in Tab. 7.
Despite being trained for only 700K steps, which is equivalent in training epochs to 350K steps with a batch size of 256 on the same dataset, LiT, using pure linear attention, achieves an impressive FID of 3.69, comparable to DiT trained for 3M steps, reducing the training steps by . Additionally, LiT outperforms several strong baselines. These results demonstrate the effectiveness of our proposed cost-efficient training strategy on higher-resolution datasets.
@@ 5.2 Text-to-Image Generation
512px resolution.
Text-to-image generation is important in generative models for commercial use. Following PixArt- [8], we add cross-attention module to our LiT-XL/2 to support text embedding input. We set the linear attention heads to 2 and the DWC kernel size to 5. Following PixArt- [7], we use pre-trained SDXL [56] VAE Encoder and T5 [10] Encoder (i.e., Flan-T5-XXL) to extract image and text features, respectively. We use PixArt- as the teacher to supervise the training of LiT, with , set to 1.0, 0.05. As in Sec. 4.3, we inherit the weights from PixArt-, except for the parameters in the self-attention. Subsequently, we train LiT on an internal dataset with a learning rate of for only 45400 steps, which is significantly fewer than the multi-stage training used for PixArt-.
Fig. 7, Fig. 11, and Fig. 12 (illustrated in Sec. E of the appendix) depict the 512px images generated by LiT. Despite using linear attention in each block, along with our cost-effective training strategy, LiT can still produce exceptional, photorealistic images. The results show the generality and commercial value of our approach for cost-effective training of efficient text-to-image frameworks.
1K resolution.
We then scale the resolution to the more challenging 1K setting. Model of 512px is selected as the starting point for training, with positional embeddings (PE) Interpolation trick applied. Token length is increased from to . We use the PixArt- [7] model, designed for 1K resolution, as the teacher model. We prepare two internal datasets: one follows the PixArt- setup, and the other follows the PixArt- setup, with a caption ratio of long (sharegpt caption) and short (original captions). The training consists of multiple stages: stage 1 trains on the first dataset, and stages 2, 3, and 4 train on the second dataset. We use 32 NVIDIA V100 GPUs, with batch sizes of 64, 64, 256, and 256 for the four stages. In the final stage, no knowledge distillation method is used.
Fig. 1 shows the sampled 1K resolution images generated by LiT. Despite replacing all self-attention with the cheap linear attention, LiT is still able to generate photo-realistic images at high resolution.
5.3 On-Device Deployment
In this section, we deploy the LiT model for 1K resolution images to an edge device to validate its on-device capability. Specifically, we deploy the model trained after stage 3 of the 1K resolution experiments from Sec. 5.2 to a laptop powered by the Windows 11 operating system. Considering the GPU memory limitations of the laptop, we quantize the text encoder to 8 bits while maintaining fp16 precision during the linear attention computation. Fig. 6 shows our deployment results. The pre-trained LiT can quickly generate photo-realistic 1K resolution images in an offline setting (without network connection). These results confirm the successful implementation of LiT as an on-device diffusion model, advancing high-resolution text-to-image tasks on edge devices.
@@ 6 Conclusion and Limitation
This paper explores using linear attention to implement diffusion Transformer for text-to-image task. We reexamine attention head number, weight inheritance, and knowledge distillation. Our model can handle class-conditional image generation, as well as synthesize photorealistic images based on instructions. Ultra-high resolution scenarios (e.g., 4K resolution) for image generation remains a limitation, which we leave in future work.
@@ Appendix C Detailed Latency and Theoretical GMACs
We use one NVIDIA V100 GPU to evaluate the latency and theoretical GMACs of the DiT-S/B models with different numbers of heads. The task was to generate 256256 resolution images with a batch size of 8, following the experimental setup of Fig. 5 (except for the GPU type), and the results are shown in Fig. 10.
@@
Appendix D Detailed Results on Class-Conditional Image Generation
@@
Appendix E More Results on Text-to-Image Generation
We provide more text-to-image results in Fig. 11 and Fig. 12. As shown, LiT can accurately generate 512px photo-realistic images in various styles, themes, and content, whether the human instructions are simple or complicated. These results demonstrate that LiT effectively learns useful knowledge from the teacher model while maintaining exceptional computational efficiency, validating the effectiveness of our proposed cost-effective training strategy.
Appendix F Code in PyTorch
PyTorch-like code of the linear attention used in our LiT for text-to-image generation is provided in Alg. 1. As shown, when training LiT for text-to-image generation, we use fp16 precision (except for the main part of linear attention, where fp32 is used) to prevent the loss from becoming NaN, basically following PixArt- [7]. We set the kernel size of the depthwise convolution to 5.
Appendix G Full Related Work
@@
As a computationally efficient alternative to self-attention, linear attention [38] reduces computational complexity from quadratic to linear and has been proven effective in both visual understanding domain [3, 22, 23, 24, 1] and language domain [59, 78, 61, 60, 6]. EfficientViT [3] introduces a multi-scale linear attention with hardware-efficient operations to obtain a general vision backbone. Flatten Transformer [23] introduces focused linear attention to address the deficiencies in focus ability and feature diversity of linear attention, incorporating a focused function and depthwise convolution (DWC). SLAB [22] simplifies focused linear attention by retaining only the DWC component and introduces a progressive re-parameterized batch normalization to adapt offline batch normalization [35] for achieving low inference latency. These studies have been validated on visual understanding. Meanwhile, our work refines a linear attention module tailored for image generation tasks and identifies the free lunch of using few heads.

### v2

@@ 3.3 Inheriting Weights from DiT
Linear attention shares the query, key, value, and output projection matrices with full attention. Despite differing computation paradigms, these weights can be directly transferred from DiT to LiT without modification. However, as shown in Fig. 6, it remains an open question whether this can accelerate convergence. We use a DiT-S/2 pre-trained for 600K iterations for ablation. We explore five different loading strategies, i.e., (1) , , and projection matrices; (2) , and projection matrices; (3) projection matrix; (4) projection matrix; and (5) the output projection matrix. Evaluation results are presented in Tab. 3. Compared to the baseline without self-attention weight loading, none of the explored strategies improve generative performance. This can be attributed to differences in computational paradigms: linear attention directly computes the product of key and value matrices, unlike self-attention. As a result, self-attention’s key and value weights provide limited benefit to linear attention. As shown in Tab. 3, inheriting these weights leads to greater performance degradation, with FID scores of 55.29 and 55.07.
@@
3.4 Distilling Both Predicted Noise and Variance
A pre-trained DiT-S/2/XL/2 serves as the teacher network to distill the student LiT-S/2. We use a DiT-S/2 pre-trained for 800K steps for weight initialization of LiT, following Guideline 3&4. Denote as the noised input latent in timestep , where , , are height, width and the number of channels of the input latent. denote network parameters of the teacher diffusion Transformer and the student LiT network , respectively. As denoising diffusion models, the output of DiT and LiT include both noise predictions and variances of the reverse diffusion process . DDPM [30] fixes the variance and optimizes model parameters using . Naturally, a straightforward idea is to align the student network’s noise prediction capabilities with those of the teacher through a mean squared error (MSE) objective:
where is the normalization factor. The impact of distilling the predicted noise is reported in Tab. 4. It shows that using a stronger teacher model (e.g., DiT-XL/2) improves the student model’s generation capbility (51.13 vs. 53.83, measured by FID). Moreover, increasing the weight of the distillation objective from 0.05 to 0.5 consistently boosts performance, suggesting that LiT can benefit from the supervision of teacher DiT.
@@ 3.4 Distilling Both Predicted Noise and Variance
As reported in IDDPM [53], a suitable variance can contribute to the initial steps of the diffusion process, thereby improving the log-likelihood. Accordingly, the authors set the variance learnable. Based on the finding, we propose to use the variance from the teacher diffusion model as a supervision for the student, but in a moderate way. The MSE between the reverse process variances of teacher and student can be calculated as:
Closing remarks.
So far, we have finalized the roadmap with 5 practical guidelines. We firmly answer how to convert a pre-trained DiT safely and efficiently to LiT while maintaining performance. Next, we will validate it on larger capacities (e.g., B/L/XL) and text-to-image generation.
4 Experiments
As this work focuses on safely and efficiently converting a pretrained baseline into a linear version without performance gap, we thus choose DiT [54] (for class-conditional image generation) and PixArt- [7] (for text-to-image generation) as baselines. The goal for LiT is to achieve comparable performance while maintaining an efficient training.
4.1 Class-Conditional Image Generation
@@
We evaluate LiT on ImageNet [14] benchmark. LiT-S/2, B/2, L/2, XL/2 configurations are consistent with DiT, except that head numbers in linear attention are 2/3/4/4, following Guideline 2 in Sec. 3.2. DWC kernel size is set to 5 for all model variants. Detailed configuration are provided in Sec. A of the Appendix. We train 400K steps with a batch size of 256. For LiT-XL/2, we extend training to 1.4M steps (20 of DiT’s 7M steps). We used the AdamW [48] optimizer with a learning rate of and weight decay set to 0. We use 250 sampling steps, aligning with DiT. We initialize LiT’s parameters using the pre-trained DiT. , in hybrid knowledge distillation, are set to 0.5 and 0.05. Evaluation metrics include FID-50K [28], Inception Score [66] and Precision/Recall [40].
As reported in Tab. 5, LiT trained for only 100K steps already outperforms the DiT trained for 400K steps across different evaluation metrics and model variants. With extra training up to 400K steps, the model’s performance continues to improve. In Tab. 6, we compare LiT-XL/2 trained for 1.4M steps with high-performance baselines. Despite having only of the training steps of DiT-XL/2, LiT still competes on par with DiT (2.32 vs. 2.27). Besides, LiT competes favorably with several U-Net [64] based baselines. The results demonstrate that our LiT, when combined with suitable optimization strategies, can be reliably used in image generation applications.
512512 ImageNet.
We further validate LiT-XL/2 on the 512px setting. Using the pre-trained DiT-XL/2 [54] as the teacher, we initialize LiT-XL/2 with its weights. For knowledge distillation, we set and to 1.0 and 0.05, respectively, and train LiT-XL/2 for only 700K steps ( of DiT’s [54] 3M steps). Notably, unlike DiT, which used a batch size of 256, we adopt a smaller batch size of 128, requiring twice the training steps to complete 1 epoch (i.e., processing the entire training dataset once). Aside from this, all other hyper-parameter settings were consistent with the 256256 experiments. The results are reported in Tab. 7. Despite being trained for only 700K steps, which is equivalent in training epochs to 350K steps with a batch size of 256 on the same dataset, LiT, using pure linear attention, achieves an impressive FID of 3.69, comparable to DiT trained for 3M steps, reducing the training steps by . Additionally, LiT outperforms several strong baselines. These results demonstrate the effectiveness of our proposed 5 practical guidelines on higher-resolution setup.
@@ 4.2 Text-to-Image Generation
| Model | #Params. | Single | Two | Count. | Colors | Pos. | Attri. | Overall |
| LDM [63] | 1.4B | 0.92 | 0.29 | 0.23 | 0.70 | 0.02 | 0.05 | 0.37 |
| SDv1.5 [63] | 0.9B | 0.97 | 0.38 | 0.35 | 0.76 | 0.04 | 0.06 | 0.43 |
| SDv2.1 [63] | 0.9B | 0.98 | 0.51 | 0.44 | 0.85 | 0.07 | 0.17 | 0.50 |
| LlamaGen [69] | 0.8B | 0.71 | 0.34 | 0.21 | 0.58 | 0.07 | 0.04 | 0.32 |
| PixArt- [8] | 0.6B | 0.98 | 0.50 | 0.44 | 0.80 | 0.08 | 0.07 | 0.48 |
| PixArt- [7] | 0.6B | - | - | - | - | - | - | 0.52 |
| Lumina-Next [89] | 2.0B | - | - | - | - | - | - | 0.46 |
| SEED-X [18] | 17B | 0.97 | 0.58 | 0.26 | 0.80 | 0.19 | 0.14 | 0.49 |
| Chameleon [70] | 34B | - | - | - | - | - | - | 0.39 |
| LWM [45] | 7B | 0.93 | 0.41 | 0.46 | 0.79 | 0.09 | 0.15 | 0.47 |
| LiT (512px) | 0.6B | 0.97 | 0.43 | 0.42 | 0.79 | 0.09 | 0.12 | 0.47 |
| LiT (1024px) | 0.6B | 0.98 | 0.50 | 0.40 | 0.77 | 0.11 | 0.12 | 0.48 |
Experimental setup.
Building up with PixArt- [7] as the baseline, we develop LiT-0.6B with 2 linear attention heads and a DWC kernel size of 5. We leverage the pretrained SDXL [57] VAE Encoder and T5 [11] Encoder (i.e., Flan-T5-XXL) for image and text feature extraction, respectively. In 512px setting, we use PixArt- as the teacher to supervise the training of LiT-0.6B, with , set to 1.0, 0.05. As in Sec. 3.3, we inherit the weights of LiT from PixArt-, except for the parameters in the whole attention module. Subsequently, we train LiT-0.6B on an internal dataset, with a mixture of long (sharegpt caption, ) and short (original caption, ) captions. Token length is set to . We train with a learning rate of and batch size of 848 for 106K steps, which is efficient compared to the multi-stage training used for PixArt- [8]. This may also be related to our proposed weight inheritance strategy. We use 20 sampling steps, following PixArt-. In 1K setting, we basically follow the 512px seting for training and qualitative evaluation. We choose 512px LiT-0.6B as the starting point for training, with positional embeddings interpolation applied. We use the PixArt- [7] for 1K resolution as the teacher model. Starting from the 512px LiT-0.6B, we continue training on the internal datasets for 155K steps with a batch size of 248. We select suitable LiT-0.6B variants for sampling qualitative results and laptop deployment.
Results.
Tab. 8 compare the GenEval [19] results of LiT with those of other powerful methods. Despite relying solely on linear attention, LiT competes on par with its baseline PixArt-, and surpasses several strong text-to-image generation baselines, such as Lumina-Next [89], LlamaGen [69], and SDv1.5 [63]. Fig. 5, 6, and 7 (Sec. F of the Appendix) depict the 512px images generated by LiT. LiT-0.6B can produce exceptional, photorealistic images. Fig. 2 shows the sampled 1K resolution images generated by LiT. Our LiT-0.6B is able to generate images at a higher resolution with fine details and colorful contents, with pure attention in each block. These results show the generality of our proposed practical guidelines in text-to-image frameworks.
| Model | #Params. | Attention | Laptop | Latency (1K) | Latency (2K) |
| PixArt- [7] | 0.6B | KV Compress | ✗ | 4.38s | 32.16s |
| LiT | 0.6B | Linear | ✓ | 3.93s | 14.59s |
Efficiency Analysis
We compare the latency of LiT and our baseline PixArt- [7] for generating 1K and 2K resolution images, as shown in Tab. 9. The results are averaged over 10 runs on a 40GB NVIDIA A100 GPU. While PixArt- employs key-value compression [76] for attention acceleration, LiT’s linear attention proves more efficient, reducing latency by at 1K and at 2K resolution.
On-Device Offline Deployment
We deploy our LiT-0.6B for 1K resolution to a laptop powered by the Windows 11 operating system to validate its on-device capability. Specifically, we quantize the text encoder to 8 bits while using fp16 precision for other weights. The results (two photos) are provided in Sec. F of the Appendix. LiT-0.6B generates photorealistic 1K-resolution images offline, showcasing its efficiency for edge computing.
@@ 5 Conclusion and Limitation
This paper answers the question of how to safely and efficiently convert a pre-trained DiT for (text-to-)image generation to a linear DiT without compromising model capacity. We reexamine attention head number, weight inheritance, and knowledge distillation, and propose 5 useful guidelines as the response. In future work, we will focus on scalability and high-resolution (e.g., 4K) LiT generation.
Acknowledgement
This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000.
@@ Appendix C Detailed Latency and Theoretical GMACs
We use one NVIDIA V100 GPU to evaluate the latency and theoretical GMACs of the DiT-S/B models with different numbers of heads. The task was to generate 256256 resolution images with a batch size of 8, following the experimental setup of Fig. 5 in the main paper (except for the GPU type), and the results are shown in Fig. 9.
@@
Appendix D Detailed Results on Attention Head Similarities
We take LiT-S/2 with 6 heads for visualization. Average cosine similarity among the attention maps of different heads is illustrated in Fig. 10.
Appendix E Detailed Results on Class-Conditional Image Generation
@@
Appendix F More Results on Text-to-Image Generation
We provide more text-to-image results in Fig. 13 and Fig. 14. As shown, LiT can accurately generate 512px photo-realistic images in various styles, themes, and content, whether the human instructions are simple or complicated. These results demonstrate that LiT effectively learns useful knowledge from the teacher model while maintaining exceptional computational efficiency, validating the effectiveness of our proposed cost-effective training strategy. Results of the offline laptop development are shown in Fig. 11.
Appendix G Pseudo-code
As presented in Alg. 1, we provide an example of the pseudo-code for the linear attention used in our LiT model for text-to-image generation for the laptop development. We set the kernel size of the depthwise convolution to 5.
Appendix H Full Related Work
@@
As a computationally efficient alternative to self-attention, linear attention [39] reduces computational complexity from quadratic to linear and has been proven effective in both visual understanding domain [3, 23, 24, 25, 1] and language domain [60, 80, 62, 61, 6]. EfficientViT [3] introduces a multi-scale linear attention with hardware-efficient operations to obtain a general vision backbone. Flatten Transformer [24] introduces focused linear attention to address the deficiencies in focus ability and feature diversity of linear attention, incorporating a focused function and depthwise convolution (DWC). SLAB [23] simplifies focused linear attention by retaining only the DWC component and introduces a progressive re-parameterized batch normalization to adapt offline batch normalization [36] for achieving low inference latency. These studies have been validated on visual perception tasks. Meanwhile, our work refines a linear attention module tailored for image generation tasks and identifies the free lunch of using few heads.